find the product -8x^2y(-5x^5y^4)
Answer by lenny460(1073)

Find the product:
-8x^2y(-5x^5y^4)
-8 * -5 = 40
x^2 * x^5 = x^7
y * y^4 = y^5
The Answer:
40x^7y^5